OpenStudy (anonymous):

EASY MEDAL JUST NEED FAST ANSWERS Write the equation y - 2 = 4(x + 5) in standard form. (Points : 1) 4x + y = 7 4x - y = 3 4x - y = -22

OpenStudy (jdoe0001):

that'd be a linear equation, so you'd want the standard form of a linear equation

OpenStudy (anonymous):

jode im running out of time on the quiz im on i need fast answers

OpenStudy (anonymous):

jdoe0001 can you please tell me the answer

OpenStudy (judygreeneyes):

Is standard form y = mx + b ?? Then 4x - y = -22 turns into -y = -4x -22 y = 4x + 22

OpenStudy (judygreeneyes):

4x - y = -22 is "standard form" for a linear equation. y = 4x + 22 is the "slope-intercept" form.